Angle ABC and angle DEB are vertical angles. Angle ABC = (4x + 50) degrees, and angle DEB = (2x + 60) degrees. Find x.
Answer by AdolphousC(70)

Vertical Angles are always equal, so set the two expressions equal and solve.


4x + 50 = 2x + 60 Subtract 2X from both sides
2x + 50 = 60 Subtract 50 from both sides
2x = 10 Divde both sides by 2
x = 5.


Check
4x + 50 = 4(5) + 50 = 20 + 50 = 70
2x + 60 = 2(5) + 60 = 10 + 60 = 70


Since we checked them and both angles are equal, we know x = 5